a plane flies at 1000m above a man. the angle of elevation is 60 degrees, suddenly the angle is 30 degrees and 5 seconds later the angle of elevation in 45 degrees. what is the speed of the plane.
Answer by lwsshak3(11628)

a plane flies at 1000m above a man. the angle of elevation is 60 degrees, suddenly the angle is 30 degrees and 5 seconds later the angle of elevation in 45 degrees. what is the speed of the plane.
***
let x=distance plane traveled in 5 sec
tan 45˚=1000/x
x=1000/tan45=1000/1=1000 m=1km
x=km/5s=km/5s*3600 s/hr
x=720 km/hr
What is the speed of the plane? 720 km/hr